OpenStudy (anonymous):

If sin y° = 9 divided by c. and tan y° = 9 divided by d. ,what is the value of cos y°?

OpenStudy (jdoe0001):

recall your SOH CAH TOA \(\bf {\color{brown}{ sin}}(\theta) =\cfrac{1}{csc(\theta)} \\ \quad \\ % cosine {\color{blue}{ cos}}(\theta) =\cfrac{1}{sec(\theta)} \\ \quad \\ % tangent tan(\theta) =\cfrac{{\color{brown}{ sin}}(\theta)}{{\color{blue}{ cos}}(\theta)}\) thus we know that \(\bf sin(y^o)=\cfrac{9}{c}\qquad tan(y^o)=\cfrac{9}{d}\qquad thus \\ \quad \\ tan(\theta) =\cfrac{{\color{brown}{ sin}}(\theta)}{{\color{blue}{ cos}}(\theta)}\implies \cfrac{9}{d}=\cfrac{\frac{9}{c}}{cos(y^o)}\impliedby \textit{solve for }cos(y^o)\)